OpenStudy (anonymous):

PLEASE HELP!!!!!!!! what is log base 3 six square root 243 (where 6 is inside the "v" of the square root"

OpenStudy (anonymous):

\[\log _{3} \sqrt[6]{243}\]

OpenStudy (anonymous):

\[\log_{3}\sqrt[6]{243} \] like this?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

alright lemme see what i can do

OpenStudy (anonymous):

okay, can u help me with a few others as well? i am really not understanding logs :(

OpenStudy (anonymous):

no problem. You know the rule of logs? Where logANS / logBASE = expo?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so in this case log [\sqrt[6]{243}\] / log3 = 5/6

OpenStudy (anonymous):

i don't really get it..can you draw it out?

OpenStudy (anonymous):